Write (–1, –2), y = –3x + 5 in slope intercept form


1
Expert's answer
2021-01-19T14:38:48-0500

I have to find here a equation which passes through "(-1,-2)" and parallel to "y=-3x+5" in slope - intercept form.

Here the given equation is "y=-3x+5" .........(1)

Slope of the given line (1) is "m=-3"

Equation of the line which is passes through "(x_1,y_1)" and having slope "m" is , "(y-y_1)=m(x-x_1)"

As the required line is parallel to (1), therefore the slope of the required line is "m=-3" and it passes through (-1,-2).

Therefore the required equation of the line is "(y+2)=(-3).(x+1)"

"\implies y=-3x-5"

Which is the required equation of the line and it is in the slope - intercept form and having slope "(-3)" and "y" intercept is "5" .